OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 36 y = x - 6 2 and 0 6 and 2 6 and -2 6 and 0

OpenStudy (anonymous):

Please help 😦

OpenStudy (xapproachesinfinity):

replace y with x-6 in the first equation and solve for x

OpenStudy (acxbox22):

^then simplify then factor and solve

OpenStudy (anonymous):

answer is B

OpenStudy (xapproachesinfinity):

don't give answers dear show how you get it

OpenStudy (xapproachesinfinity):

\[x^2+(x-6)2=36\\ x^2+x^2-12x+36=36\] \[2x^2-12x=0 \\ 2x(x-6)=0\] \[2x=0~~or ~~x-6=0\]

OpenStudy (anonymous):

I got 6, -7. I don't know how I went wrong or where.

OpenStudy (xapproachesinfinity):

finish up the work

OpenStudy (xapproachesinfinity):

that's incorrect

OpenStudy (anonymous):

Ahhh, I see what I did.